3. Describe verbally how to solve y=mx+c. What assumptions have you made about the value

of ?


Added NOTE:

For question 3, students should just given a brief description, in writing, of how you would solve  y=mx+c for x.  Don't solve it, just say how you would do it and answer the remainder of the question as well.


1
Expert's answer
2021-10-13T17:22:21-0400

y=mx+c is the equation of straight line where y is the dependent and x is the independent variable. m is the slope and c is the intercept cut. We can choose any value of x, corresponding we obtained the value of y. The slope which is m=(y2-y1)/(x2-x1)= contant and c is the intercept cut.

Again in order to obtain x from above equation we need to rearrange the things as y=mx+c

(y-c)/m=x

Alternatively we need to take the different value of x on y axis and the value of y on x axis